# Agent Control Protocol ACP v1.19: Admission Control for Agent Actions Technical Specification and Reference Implementation

### 25 Mar 2026 v5

@@ 12.10  Roadmap
ACP-SIGN-2.0 HYBRID stub (pkg/sign2/) v1.17 Complete
@@
Post-quantum Go implementation (Dilithium, circl) v1.20
@@
Table 17: ACP v1.19 development roadmap.
@@
We introduce a compliance runner, ACR-1.0 (ACP Compliance Runner 1.0), that validates ACP-RISK-2.0 implementations against sequence-based test cases. The runner operates in two modes: library mode (direct call to pkg/risk, no network, deterministic) and HTTP mode (external server validation for interoperability). All 73 signed single-shot conformance vectors plus 5 sequence scenarios are validated on every commit.
@@ 13.1  Sequence-Based Test Vectors
Listing 1: ACR-1.0 execution contract in library mode.
@@ 14  End-to-End Verifiability
ACP v1.19 enables end-to-end verifiability by aligning formal specification, test vectors, and runtime validation through a shared request abstraction. Invariants proven in TLA+ are instantiated as concrete test vectors and validated empirically by the compliance runner.
@@
What is formally verified in TLA+ holds at runtime. Each TLA+ invariant is instantiated as a concrete test vector and validated empirically by the compliance runner. This is the connection across the three layers of the v1.19 verifiability stack: formal $\to$ data $\to$ runtime.
@@ 15.2  Additional Requirements for L3 Conformance
  • Deterministic risk engine implementing ACP-RISK-2.0 (F_res, F_ctx, F_hist, F_anom, Cooldown).

### 30 Apr 2026 v10

@@ 13.9  Summary
While ACP enforces admissibility at execution time through stateful evaluation, this guarantee relies on an implicit assumption: that the system continues to generate actions that meaningfully interact with the admissibility boundary. In practice, this assumption may fail.
@@ 13.10  Roadmap
ACP-SIGN-2.0 HYBRID stub (pkg/sign2/) v1.17 (superseded by v1.20)
ML-DSA-65 (Dilithium mode3) real implementation via CIRCL (pkg/sign2/) v1.20 Complete
@@
Token replay adversarial evaluation (Experiment 5, F_anom Rule 3 escalation) v1.20 Complete
Deployment Considerations section (state backend selection, cross-agent boundary, policy tuning) v1.20 Complete
[…]
Post-quantum Go implementation (Dilithium, circl) v1.20 Complete
NullQuerier + StatelessEngine (stateless comparison baseline) v1.21 Complete
Experiment 6: Stateless vs. ACP (0.4% vs. 100% autonomous execution rate) v1.21 Complete
Experiment 7: State-Mixing Vulnerability (cross-context Rule 1 contamination) v1.21 Complete
ACP-RISK-3.0: context-scoped Rule 1 (CountPattern replaces CountRequests) v1.22 Complete
Experiment 8: Context-Scoped Anomaly Fix (state-mixing eliminated, enforcement preserved) v1.22 Complete
Experiment 9: Deviation Collapse and Restoration (BAR metric, counterfactual evaluation) v1.23 Complete
Boundary Activation Monitoring (pkg/barmonitor, $\Delta\mathrm{BAR}$ trend detection) v1.24 Complete
EvaluateCounterfactual API (pkg/risk, additive mutations, 3 built-in factories) v1.24 Complete
Experiment 9 Phase D: $\Delta\mathrm{BAR}$ early-warning validation (drift simulation, 5 batches) v1.25 Complete
TLA+ invariants: FailureConditionPreservation, NoDegenerateAdmissibility v1.25 Complete
POST /acp/v1/counterfactual HTTP endpoint (structural + behavioral mutations) v1.25 Complete
Formal adversary model $\mathcal{A}=(K,S,B)$ + experiment taxonomy table v1.26 Complete
Experiment 11: Threshold sensitivity analysis (5 configs, monotonicity verified) v1.26 Complete
Detection Guarantees: Proposition 1, binomial detection probability v1.26 Complete
AgentSpec functional comparison (5 dimensions, Exp 10 governance collapse scenario) v1.26 Complete
§13.8: IPI scope + AgentDojo/InjecAgent benchmarks v1.27 Complete
§4.5: LangGraph/Ollama integration pseudocode v1.27 Complete
Experiment 12: Multi-tool IPI admission control, stateful Fanom persistence v1.27 Complete
Experiment 13: Bounded coordination window ($\mathrm{CW}=2N$, exact linearity, evaluate-then-mutate) v1.28 Complete
Experiment 14: OPA vs. ACP capability comparison (stateless vs. stateful, 3 scenarios) v1.29 Complete
Governance series integration: Paper 3/4 cited in §3.5; series updated to 7 papers (P3/4 consolidated, P6 added) v1.30 Complete
ACP-LEDGER-1.0 (tamper-evident hash-chained audit log for state backend integrity) Future work
@@
Table 26: ACP v1.30 development roadmap.
@@
We introduce a compliance runner, ACR-1.0 (ACP Compliance Runner 1.0), that validates ACP-RISK-2.0 and ACP-RISK-3.0 implementations against sequence-based test cases. The runner operates in two modes: library mode (direct call to pkg/risk, no network, deterministic) and HTTP mode (external server validation for interoperability). All 73 signed single-shot conformance vectors plus 5 sequence scenarios are validated on every commit.
@@ 14.1  Sequence-Based Test Vectors
Listing 3: ACR-1.0 execution contract in library mode.
@@ 15  End-to-End Verifiability
ACP v1.25 enables end-to-end verifiability by aligning formal specification, test vectors, and runtime validation through a shared request abstraction. Invariants proven in TLA+ are instantiated as concrete test vectors and validated empirically by the compliance runner.
@@
What is model-checked in TLA+ over a bounded state space is instantiated as concrete test vectors and validated empirically by the compliance runner. This is the connection across the three layers of the v1.25 verifiability stack: formal $\to$ data $\to$ runtime. The TLA+ model verifies selected safety and liveness properties; the compliance runner validates the reference implementation against those properties.
Structural failure condition preservation (v1.25). Two invariants are added to the extended TLA+ model (ACP_Extended.tla) to establish necessary structural conditions for failure condition preservation (Section 17).
FailureConditionPreservation asserts that the evaluation action space contains at least one (capability, resource) pair that the engine evaluates as DENIED under empty ledger state and default policy. Formally: $\exists\,\mathit{cap}\in\text{Capabilities},\,\mathit{res}\in\text{Resources}:\mathit{Decide}(\mathit{ComputeRiskWithAnom}(\mathit{cap},\mathit{res},0))=\texttt{DENIED}$. With the model constants (Capabilities = {read, financial, admin}, Resources = {public, sensitive}), the pair (admin, sensitive) yields $\mathit{RS}=75\geq 70\to\texttt{DENIED}$.
NoDegenerateAdmissibility asserts that high-risk capability/resource combinations are never APPROVED: for all ledger entries where capability $\in$ {admin, financial} and resource == sensitive, the decision is not APPROVED. This holds because $\mathit{ComputeRisk}(\texttt{admin},\texttt{sensitive})=75$ and $\mathit{ComputeRisk}(\texttt{financial},\texttt{sensitive})=50$, both strictly above the APPROVED threshold of 39.
Both invariants are verified by TLC with zero violations across all reachable states. These establish necessary structural conditions: they verify that the model’s action space preserves enforcement capacity. Sufficient conditions for runtime failure condition preservation under arbitrary upstream transformations require pipeline-level verification, as validated empirically in Experiment 9 Phase D.
BARMonitorLiveness (stated, not submitted to TLC): if the last $N$ decisions for an agent are all APPROVED, BAR-Monitor eventually emits an alert ($\mathrm{BAR}_{N}<\theta$). Adding monitor state to TLC increases the state space by a factor of $2^{N}$; for $N=20$ (Phase D batch size) this is not tractable. The property is validated empirically in Phase D (Table 30).
@@ 16.2  Additional Requirements for L3 Conformance
  • Deterministic risk engine implementing ACP-RISK-3.0 (F_res, F_ctx, F_hist, F_anom, Cooldown), which supersedes ACP-RISK-2.0 while preserving API- and policy-level compatibility.
@@
16.3  Counterfactual Evaluation Endpoint
The reference server exposes a counterfactual evaluation endpoint (POST /acp/v1/counterfactual) that allows external tooling to verify that a deployment retains structural enforcement capacity. A caller submits a base request and a list of mutations; the server evaluates each mutation independently against the risk engine and returns the per-mutation decision, risk score, and aggregate BAR.
The endpoint supports structural mutations (elevated capability and resource class) and behavioral mutations (injected context and history flags). Temporal mutations (requiring pre-loaded ledger state via LedgerSetup) are not supported via HTTP; they are available through the Go library API (pkg/risk.TemporalMutation).
This endpoint is intended for conformance testing, operational health checking, and third-party verification that a deployed ACP instance can produce DENIED decisions—a necessary condition for meaningful governance.